Lets say you want to float a section of copper wire, which is 7.65 cm long and 2.34 mm in diameter within the magnetic fie
Answer #1

Length of the copper wire L=7.65 cm= 7.65*10-2 m

Diameter of the wire=2.34 mm

Radius r=2.34/2=1.17*10-3 m

Angle made by the magnetic field with horizontal \theta =31o

Magnetic field of earth=3.5*10-5T

Thus, horizontal component of the magnetic field, BV=B sin\theta=3.5*10-5*sin31=1.8*10-5T

5) The free body diagram of the current carrying wire is shown below.

Bc = po I 2r I BH N. 31° Bu To6) To suspend or float a section of copper wire, the magnetic field due to the current carrying wire shorld be equal to the vertical component of the earth's magnetic field as shown above.

Thus, \mu oI/2r=1.8*10-5

where \mu o= permeability of free space=4\pi*10-7N/A2

Using r=1.17*10-3 m(given),

Current through the wire I=1.8*10-5*3.34*10-3/(4\pi*10-7)=0.048A

7) According to Faraday's law emf e=-d\phi/dt, where \phi =magnetic flux=B.dA

For the wire, dA=2\pirL

Thus, e=-d/dt( \mu oI/2r*2\pirL)=-d/dt( \pi \muoI/L)

This will change the current in the loop, in which emf changes according to the Lenz law in a direction opposite to the source producing it, thus increasing or decreasing the current.

The vertical position remains unchanged since the magnetic field and hence the current is independent of the vertical length L.